# Supervising 3D Talking Head Avatars with Analysis-by-Audio-Synthesis

### v1

@@ 1 Introduction
Seminal approaches such as VOCA [12] or Karras et al. [35], and others [19, 14, 48] are deterministic regressors, that predict 3D vertices from the input audio. Despite achieving good lip-sync, these methods produce facial animations without emotion [12, 19] or that are unnaturally static [48, 14]. This is in part due to the lack of large-scale and expression-rich training data and in part because a deterministic regressor cannot accurately produce animations of facial motions that only weakly correlate with the audio (such as eyebrow motion, eyeblinks, etc.). Hence, in order to generate animations that appear lively and believable, a good 3D speech-driven system must model the problem in a non-deterministic manner so that it can capture the many-to-many mapping between speech and facial animation.
Consequently, recent methods model the problem of 3D speech-driven animation stochastically using relatively small datasets of high-quality 3D scans [56, 71, 59, 65]; the small training size limits diversity and expressiveness. Other methods leverage lower-quality pseudo ground-truth (GT) extracted from larger-scale video datasets [78, 60]. In pursuing diversity, however, stochastic methods vary the lip motions in ways that deviate from the audio signal.
The fundamental issue is that lip motions, for the same audio, vary with expression. We seek a method that generates diverse expressions but is faithful to the audio. So we ask the question: “What does it mean to be faithful to the audio?” This leads us to our key observation; that is, we seek 3D lip motions that actually reproduce the input audio. We formalize this notion with a new form of supervision for 3D speech-driven avatars using a novel mesh-to-speech model; see Fig. 1. Inspired by recent silent-video-to-speech (SVTS) methods [8, 37, 36], we propose, to our knowledge, the first mesh-to-speech model. Our mesh-to-speech model regresses the speech audio (or a representation thereof) solely from the animated 3D face. In other words, every generated animation now also produces a sound. Our rationale is simple: the more similar the produced sound is to the input speech, the more plausible the animation. Furthermore, failure to produce the correct sound should provide a good supervision signal.
@@
In summary, our contributions are: (1) A novel mesh-to-speech model that regresses sound from 3D facial animations. (2) A new form of cross-modal audio-based self-supervision via comparison of the representations of the input speech and the animation-produced output speech. (3) A 3D speech-driven animation method based on diffusion, capable of generating a variety of facial expressions while maintaining accurate lip-sync. The code and models will be made publicly available at https://thunder.is.tue.mpg.de/ .
@@ 2.1 Speech-driven 3D animation
The field of speech-driven 3D facial animation has made significant progress in the last two decades [4, 17, 18, 72, 62, 9]. Here we focus on the recent deep-learning based line of work [50, 49, 35, 63, 79, 12, 56, 19, 71, 65, 59, 78, 60, 75, 48, 45, 46].
@@
Deterministic neural methods. Karras et al. [35] were the first to utilize deep learning by training a temporal convolutional network (TCN) to predict 3D face vertices. VOCA [12] employs a pre-trained automatic speech recognition (ASR) network to regress face vertex offsets from audio, achieving good lip-sync for multiple speaking styles. Many follow-up methods use a similar paradigm [19, 66, 23, 48, 14, 45, 46, 73, 2].
@@
Stochastic neural methods. The first model to approach the problem stochastically was MeshTalk [56]. MeshTalk makes use of a pretrained discretized facial motion prior along with an explicit supervision mechanism that supports generation of motions that only have a weak correlation with the audio (eyebrow motion, etc.). CodeTalker [71] upgrades the FaceFormer architecture by predicting codebook classes of a pretrained VQ-VAE. Both MeshTalk and CodeTalker employ autoregressive predicition and one can therefore sample the distribution of the next token to generate variety.
@@
Thanks to the tremendous success of Diffusion Models in the image domain [26, 58] they are now widely used for 3D human body animation [64, 5], as well as speech-driven animation [59, 65, 60, 78]. The diffusion methods also employ transformer-based audio feature extractors [3, 28] to condition the denoiser. FaceDiffuser [59] employs a GRU-based denoiser, 3DiFACE [65] uses a TCN-based denoiser, and DiffPoseTalk [60] and Media2Face [78] employ a transformer decoder architecture [68] passing the audio condition to the denoiser via cross-attention.
@@
Controlling the output animation. The task of controlling or editing the output animation is of great utility in production as manually editing animations is a laborious process. Many methods use a simple one-hot encoding of training subjects, which is mapped to a “style embedding” with a learnable layer. This embedding conditions the decoder to match the speaking style of the corresponding training subject [12, 19, 71, 66]. This paradigm has been extended to control the emotion and emotion intensity [14, 48]. DiffPoseTalk [60] uses contrastive learning to produce a style vector from a reference animation, lifting the subject-ID conditioning limitation. Media2Face [78] employs CLIP features [53], which can be extracted from an image, or text prompts to condition the denoising process. 3DiFACE [65] demonstrates that the output of a diffusion model can be guided by a sparse set of keyframes. In this paper, instead of focusing on controlling the animation, we present a general method to improve lip-sync quality, which can be used in conjunction with the above control mechanisms.
@@
Datasets. Most methods [35, 12, 19, 71, 66] have made use of high quality 3D scans synchronized with audio, such VOCASET [12], BIWI [20] or Multiface [70]. While these datasets are of high quality, they are expensive to acquire and hence are limited in number of subjects, and richness of facial expressions. The dataset design is often a limiting factor, as the samples are usually sentences taken out of context and hence the subjects do not express themselves in the most natural ways. As a consequence, even the stochastic methods that leverage these datasets [71, 56, 59], suffer from lack of natural diversity of expressions.
@@
Thanks to the significant improvement of in-the-wild face reconstruction systems [13, 80, 21, 77, 55] that can now be utilized off-the-shelf, recent methods have turned to using pseudo-ground-truth (pGT) reconstructions from videos [14, 48, 60, 78]. While the pGT does not reach the quality of 3D scans, video datasets are numerous, allowing more data to be acquired, which benefits data-hungry stochastic models.
@@ 2.2 Silent-video-to-speech (SVTS)
To the best of our knowledge, the prediction of voice from a sequence of 3D facial shapes has never been explored. However, recent years have seen tremendous progress on the task of speech audio prediction from silent videos. Early methods [11, 47, 27, 36, 44, 74, 69, 67] focus on small in-the-lab datasets with predefined scripts and limited number of speakers (GRID [10], TCD-TIMIT [24]) or script-unconstrained but single-speaker models [51].
STVS [16] was the first method to produce intelligible audio on large-scale in-the-wild datasets such as LRS2 and LRS3 [1] by regressing spectrograms from mouth crops. The spectrograms are converted to the final waveform with a pretrained vocoder[8]. Follow-up methods improve the prediction quality [37] by adding a surrogate ASR loss or predicting features from an ASR model [8]
The most recent architectures are based on diffusion, producing audio often indistinguishable from real speech [7, 76]. These, however, are not suitable for a self-supervised reconstruction loop, due to the iterative nature of diffusion models.
Hence, thanks to its simple, yet efficient feed-forward design and a relatively straightforward architecture and training, we base our mesh-to-speech model on Choi et al.[8].
@@ 2.3 Audio cycle consistency
Cycle consistency losses on audio have been applied before, for tasks like voice conversion [32, 33, 34] and disentangled representation learning [31, 6]. Many different self-supervised audio learning methods exist (see survey [41]). However, cross-modal cycle consistency between audio and 3D facial motion has, to our knowledge, never been proposed.
@@ 3.1 Face model
THUNDER uses the FLAME face model [40], which is a statistical 3D morphable model (3DMM). FLAME provides a compact representation of facial shapes and expressions and is defined as a function:
@@ 3.2 Speech feature extraction
Following previous work on speech-driven facial animation [19, 71, 66, 14, 78] we use Wav2Vec2.0 [3] as our audio feature extractor. It consists of a temporal convolution network (TCN) that extracts audio features at 50Hz and a transformer encoder that processes these features. Similarly to previous methods, we resample the TCN feature to 25Hz and feed it to the transformer to obtain the final speech feature. Formally, we define this as: $\mathcal{A}(\mathbf{w})\rightarrow\mathbf{a}^{1:T},$ where $\mathcal{A}$ denotes the Wav2Vec 2.0 network, $\mathbf{w}$ is the input waveform, and $\mathbf{a}^{1:T}\in\mathbb{R}^{T\times d_{s}}$, with $T$ denoting the number of frames at 25Hz and $d_{s}=768$ is the feature dimension.
@@ 3.3 Speaker Embedding
A speaker embedding is a numerical representation of a speaker’s unique vocal characteristics encoded into a vector. This vector captures features such as tone, pitch, speaking style, and other vocal attributes, enabling the differentiation and recognition of individual speakers. Among other applications, speaker embeddings are used in the latest SVTS models [16, 37, 8]. Following Choi et al. [8], we leverage an off-the-shelf speaker embedding extractor [29]. We define the extractor as: $\mathcal{S}(\mathbf{w})\rightarrow\boldsymbol{s},$ with $\mathcal{S}$ and $\boldsymbol{s}$ denoting embedding extractor and the embedding vector, respectively.
@@ 3.4 Speech Units
The term speech units refers to discrete linguistic units identified through a self-supervised learning process. Like Choi et al. [8], we leverage speech units predicted by a pretrained HuBERT [28] model and its associated feature clustering model. First, features are extracted and subsequently clustered into $n_{\boldsymbol{U}}$ units. We define speech units as one-hot class vectors $\boldsymbol{U}\in[c_{1},c_{2},\dots,c_{n_{\boldsymbol{U}}}]$ where only one of $c_{i},i=1\dots c_{n_{\boldsymbol{U}}}$ is $1$ and the rest are 00. Speech units are predicted at 50Hz.
@@ 4.1 Mesh-To-Speech
Regressing audio from 3D facial motion has, to our knowledge, not been done before. To narrow down the architecture search space, we take inspiration from a recent SVTS model, namely Choi et al. [8].
Architecture. The architecture consists of an input feature encoder, a conformer sequence encoder and two prediction heads, a speech unit classifier, and a mel spectrogram regressor. These two outputs contain enough information that a pretrained off-the-shelf vocoder [8] can turn them into output audio. Since the input is not a video, but a 3D animation, we replace the original lip video ResNet encoder that Choi et al. [8] use with an MLP that takes 3D lip vertex coordinates as the input. We keep the rest of the architecture the same. The mesh-to-speech architecture is depicted in Fig. 2 and can be written
@@
Supervision. We follow the same supervision scheme as Choi et al. [8]. The loss consists of two terms. The first term is an L1 loss between the input and output mel specrograms:
@@ 4.2 Speech-Driven Facial Animation Diffusion
Architecture. The core of THUNDER consists of a denoiser in the form of a transformer decoder [68]. Similar to other contemporary 3D animation methods (e.g. Zhao et al. [78] and Sun et al. [60]), THUNDER processes a noisy sequence of expression parameters $\boldsymbol{\tilde{x}}_{(d)}^{1:T}$, where $d$ represents the diffusion timestep. These parameters serve as the query inputs to the transformer. The keys and values for the transformer are comprised of the input audio features $\mathbf{a}^{1:T}$ and a diffusion step embedding $\mathcal{T}(d)$, concatenated along the temporal dimension. The function $\mathcal{T}$ is defined as a sinusoidal timestep function similar to Ho et al. [26]. We do not incorporate additional animation-controlling conditions (such as style vectors [12, 19, 71], emotion vectors [14, 48] or CLIP embeddings [78]) as this is not the goal of this work. However, the design can easily be augmented with additional input conditions. All inputs are projected into the transformer’s feature dimension, $f=128$, using dedicated learnable affine layers. Positional encoding within the transformer’s self-attention utilizes the ALiBi mechanism [52]. The cross-attention layers use a diagonal binary mask to facilitate the attention mechanism, and no additional positional encoding. The model predicts the reconstructed expression parameters $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$. The transformation can be formally described by the following equation:
@@
where $\alpha_{d}$ represents the proportion of the original signal’s variance preserved at diffusion timestep $d$, and $\epsilon$ is drawn from a Gaussian distribution with predefined $\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}$. This approach incrementally corrupts the clean parameters, enabling the model to learn to denoise varying levels of noise. The model is supervised with several loss terms - reconstruction and velocity losses in expression space and 3D vertex space, and also the mesh-to-speech loss (Eq. (5)). The reconstruction loss is given as:
@@
The velocity loss is also an MSE error but computed on the velocities of the individual components:
@@
$\mathcal{L}_{total}=w_{mts}\mathcal{L}_{m2s}+\mathcal{L}_{rec}+\mathcal{L}_{vel},$ (8)
with $w_{mts}=1$. The model is trained using classifier-free guidance [25]. The audio condition is randomly dropped and replaced with a learnable vector with 20% probability.
Inference. At inference time, we initialize the first step by sampling the Gaussian noise: $\boldsymbol{\widehat{x}}_{(D)}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$. We then proceed to denoise for $D$ diffusion steps. In each step, we first employ the denoiser $\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\mathbf{a}^{1:T})$ to obtain $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$. Then, we compute $\boldsymbol{\tilde{x}}_{(d-1)}$ by adding the expected amount of noise back to obtain $\boldsymbol{\tilde{x}}_{(d-1)}$. Since THUNDER is trained with classifier-free guidance, it is possible to trade-off fidelity of the lip-sync for increased diversity by combining $s_{a}\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\mathbf{a}^{1:T})+(1-s_{a})\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\O)$ during the denoising process. However, we find our results are diverse enough without this and so we effectively set $s_{a}=0$.

### v2

@@ 1 Introduction
Seminal approaches such as VOCA [cudeiro2019capture] or Karras et al. [karras2017audio], and others [fan2022faceformer, danecek2023emote, Peng2023_EmoTalk] are deterministic regressors, that predict 3D vertices from the input audio. Despite achieving good lip-sync, these methods produce facial animations without emotion [cudeiro2019capture, fan2022faceformer] or that are unnaturally static [Peng2023_EmoTalk, danecek2023emote]. This is in part due to the lack of large-scale and expression-rich training data and in part because a deterministic regressor cannot accurately produce animations of facial motions that only weakly correlate with the audio (such as eyebrow motion, eyeblinks, etc.). Hence, in order to generate animations that appear lively and believable, a good 3D speech-driven system must model the problem in a non-deterministic manner so that it can capture the many-to-many mapping between speech and facial animation.
Consequently, recent methods model the problem of 3D speech-driven animation stochastically using relatively small datasets of high-quality 3D scans [richard2021meshtalk, xing2023codetalker, stan2023facediffuser, thambiraja20233diface]; the small training size limits diversity and expressiveness. Other methods leverage lower-quality pseudo ground-truth (GT) extracted from larger-scale video datasets [zhao2024media2face, sun2024diffposetalk]. In pursuing diversity, however, stochastic methods vary the lip motions in ways that deviate from the audio signal.
The fundamental issue is that lip motions, for the same audio, vary with expression. We seek a method that generates diverse expressions but is faithful to the audio. So we ask the question: “What does it mean to be faithful to the audio?” This leads us to our key observation; that is, we seek 3D lip motions that actually reproduce the input audio. We formalize this notion with a new form of supervision for 3D speech-driven avatars using a novel mesh-to-speech model; see Fig. 1. Inspired by recent silent-video-to-speech (SVTS) methods [choi2023lip2speechunit, kim2023lip, kim2021lipvca], we propose, to our knowledge, the first mesh-to-speech model. Our mesh-to-speech model regresses the speech audio (or a representation thereof) solely from the animated 3D face. In other words, every generated animation now also produces a sound. Our rationale is simple: the more similar the produced sound is to the input speech, the more plausible the animation. Furthermore, failure to produce the correct sound should provide a good supervision signal.
@@
In summary, our contributions are: (1) A novel mesh-to-speech model that regresses sound from 3D facial animations. (2) A new form of cross-modal audio-based self-supervision via comparison of the representations of the input speech and the animation-produced output speech. (3) A 3D speech-driven animation method based on diffusion, capable of generating a variety of facial expressions while maintaining accurate lip-sync. The code and models will be made publicly available .
@@ 2.1 Speech-driven 3D animation
The field of speech-driven 3D facial animation has made significant progress in the last two decades [cao2005expressive, edwards2016jali, edwards2020jalicyberpunk, xu2013practical, taylor2012dynamic, cohen2001animated]. Here we focus on the recent deep-learning based line of work [pham2017end, pham2017speech, karras2017audio, taylor2017generalized, zhou2018visemenet, cudeiro2019capture, richard2021meshtalk, fan2022faceformer, xing2023codetalker, danecek2023emote, thambiraja20233diface, stan2023facediffuser, zhao2024media2face, sun2024diffposetalk, yang2024probabilistic, Peng2023_EmoTalk, nocentini2024scantalk, aneja2024facetalk, aneja2024gaussianspeech, nocentini2024emovoca, enugi2024enhancing, fan2024unitalker].
@@
Deterministic neural methods. Karras et al. [karras2017audio] were the first to utilize deep learning by training a temporal convolutional network (TCN) to predict 3D face vertices. VOCA [cudeiro2019capture] employs a pre-trained automatic speech recognition (ASR) network to regress face vertex offsets from audio, achieving good lip-sync for multiple speaking styles. Many follow-up methods use a similar paradigm [fan2022faceformer, bala2022imitator, FaceXHuBERT_Haque_ICMI23, Peng2023_EmoTalk, danecek2023emote, peng2023selftalk, nocentini2024scantalk, nocentini2024emovoca, xu2024kmtalk, fu2024mimic, aneja2024gaussianspeech, enugi2024enhancing, fan2024unitalker].
@@
Stochastic neural methods. The first model to approach the problem stochastically was MeshTalk [richard2021meshtalk]. MeshTalk makes use of a pretrained discretized facial motion prior along with an explicit supervision mechanism that supports generation of motions that only have a weak correlation with the audio (eyebrow motion, etc.). CodeTalker [xing2023codetalker] upgrades the FaceFormer architecture by predicting codebook classes of a pretrained VQ-VAE. Both MeshTalk and CodeTalker employ autoregressive predicition and one can therefore sample the distribution of the next token to generate variety.
@@
Thanks to the tremendous success of Diffusion Models in the image domain [ho2020denoising, rombach2022stable] they are now widely used for 3D human body animation [tevet2023hmd, chen2023mld], as well as speech-driven animation [stan2023facediffuser, thambiraja20233diface, sun2024diffposetalk, zhao2024media2face]. The diffusion methods also employ transformer-based audio feature extractors [baevski2020wav2vec, hsu2021hubert] to condition the denoiser. FaceDiffuser [stan2023facediffuser] employs a GRU-based denoiser, 3DiFACE [thambiraja20233diface] uses a TCN-based denoiser, and DiffPoseTalk [sun2024diffposetalk], Media2Face [zhao2024media2face] and FaceTalk [aneja2024facetalk] employ a transformer decoder architecture [vaswani2017attention] passing the audio condition to the denoiser via cross-attention.
@@
Controlling the output animation. The task of controlling or editing the output animation is of great utility in production as manually editing animations is a laborious process. Many methods use a simple one-hot encoding of training subjects, which is mapped to a “style embedding” with a learnable layer. This embedding conditions the decoder to match the speaking style of the corresponding training subject [cudeiro2019capture, fan2022faceformer, xing2023codetalker, bala2022imitator]. This paradigm has been extended to control the emotion and emotion intensity [danecek2023emote, Peng2023_EmoTalk]. DiffPoseTalk [sun2024diffposetalk] uses contrastive learning to produce a style vector from a reference animation, lifting the subject-ID conditioning limitation. Media2Face [zhao2024media2face] employs CLIP features [Radford2021clip], which can be extracted from an image, or text prompts to condition the denoising process. 3DiFACE [thambiraja20233diface] demonstrates that the output of a diffusion model can be guided by a sparse set of keyframes. In this paper, instead of focusing on controlling the animation, we present a general method to improve lip-sync quality, which can be used in conjunction with the above control mechanisms.
@@
Datasets. Most methods [karras2017audio, cudeiro2019capture, fan2022faceformer, xing2023codetalker, bala2022imitator] have made use of high quality 3D scans synchronized with audio, such VOCASET [cudeiro2019capture], BIWI [eth_biwi_00760] or Multiface [wuu2022multiface]. While these datasets are of high quality, they are expensive to acquire and hence are limited in number of subjects, and richness of facial expressions. The dataset design is often a limiting factor, as the samples are usually sentences taken out of context and hence the subjects do not express themselves in the most natural ways. As a consequence, even the stochastic methods that leverage these datasets [xing2023codetalker, richard2021meshtalk, stan2023facediffuser], suffer from lack of natural diversity of expressions.
@@
Thanks to the significant improvement of in-the-wild face reconstruction systems [EMOCA:CVPR:2021, MICA:ECCV2022, filntisis2022visual, zhang2023tokenface, Retsinas_2024_CVPR_Smirk] that can now be utilized off-the-shelf, recent methods have turned to using pseudo-ground-truth (pGT) reconstructions from videos [danecek2023emote, Peng2023_EmoTalk, sun2024diffposetalk, zhao2024media2face]. While the pGT does not reach the quality of 3D scans, video datasets are numerous, allowing more data to be acquired, which benefits data-hungry stochastic models.
@@ 2.2 Silent-video-to-speech (SVTS)
To the best of our knowledge, the prediction of voice from a sequence of 3D facial shapes has never been explored. However, recent years have seen tremendous progress on the task of speech audio prediction from silent videos. Early methods [cornu2015speechfromfeaturs, Oneata2021speaker, hong2021speechrec, kim2021lipvca, Michelsanti2020vocoderbased, Yadav2021speech, Vougioukas2019videodriven, um2023facetron] focus on small in-the-lab datasets with predefined scripts and limited number of speakers (GRID [cooke2016grid], TCD-TIMIT [Harte2015tcdtimit]) or script-unconstrained but single-speaker models [prajwal2020cvrplearning].
STVS [carrillo2022svts] was the first method to produce intelligible audio on large-scale in-the-wild datasets such as LRS2 and LRS3 [afouras2018lrs3] by regressing spectrograms from mouth crops. The spectrograms are converted to the final waveform with a pretrained vocoder[choi2023lip2speechunit]. Follow-up methods improve the prediction quality [kim2023lip] by adding a surrogate ASR loss or predicting features from an ASR model [choi2023lip2speechunit]
The most recent architectures are based on diffusion, producing audio often indistinguishable from real speech [choi2023diffv2s, yemini2023lipvoicer]. These, however, are not suitable for a self-supervised reconstruction loop, due to the iterative nature of diffusion models.
Hence, thanks to its simple, yet efficient feed-forward design and a relatively straightforward architecture and training, we base our mesh-to-speech model on Choi et al.[choi2023lip2speechunit].
@@ 2.3 Audio cycle consistency
Cycle consistency losses on audio have been applied before, for tasks like voice conversion [Kaneko2020CycleGAN-VC, Kaneko2020CycleGAN-VC2, Kaneko2020CycleGAN-VC3] and disentangled representation learning [ji2021evp, Chhatre2024amuse]. Many different self-supervised audio learning methods exist (see survey [liu2022audioself]). However, cross-modal cycle consistency between audio and 3D facial motion has, to our knowledge, never been proposed.
@@ 3.1 Face model
THUNDER uses the FLAME face model [FLAME:SiggraphAsia2017], which is a statistical 3D morphable model (3DMM). FLAME provides a compact representation of facial shapes and expressions and is defined as a function:
@@ 3.2 Speech feature extraction
Following previous work on speech-driven facial animation [fan2022faceformer, xing2023codetalker, bala2022imitator, danecek2023emote, zhao2024media2face] we use Wav2Vec2.0 [baevski2020wav2vec] as our audio feature extractor. It consists of a temporal convolution network (TCN) that extracts audio features at 50Hz and a transformer encoder that processes these features. Similarly to previous methods, we resample the TCN feature to 25Hz and feed it to the transformer to obtain the final speech feature. Formally, we define this as: $\mathcal{A}(\mathbf{w})\rightarrow\mathbf{a}^{1:T},$ where $\mathcal{A}$ denotes the Wav2Vec 2.0 network, $\mathbf{w}$ is the input waveform, and $\mathbf{a}^{1:T}\in\mathbb{R}^{T\times d_{s}}$, with $T$ denoting the number of frames at 25Hz and $d_{s}=768$ is the feature dimension.
@@ 3.3 Speaker Embedding
A speaker embedding is a numerical representation of a speaker’s unique vocal characteristics encoded into a vector. This vector captures features such as tone, pitch, speaking style, and other vocal attributes, enabling the differentiation and recognition of individual speakers. Among other applications, speaker embeddings are used in the latest SVTS models [carrillo2022svts, kim2023lip, choi2023lip2speechunit]. Following Choi et al. [choi2023lip2speechunit], we leverage an off-the-shelf speaker embedding extractor [RealTimeVoiceCloning]. We define the extractor as: $\mathcal{S}(\mathbf{w})\rightarrow\boldsymbol{s},$ with $\mathcal{S}$ and $\boldsymbol{s}$ denoting embedding extractor and the embedding vector, respectively.
@@ 3.4 Speech Units
The term speech units refers to discrete linguistic units identified through a self-supervised learning process. Like Choi et al. [choi2023lip2speechunit], we leverage speech units predicted by a pretrained HuBERT [hsu2021hubert] model and its associated feature clustering model. First, features are extracted and subsequently clustered into $n_{\boldsymbol{U}}$ units. We define speech units as one-hot class vectors $\boldsymbol{U}\in[c_{1},c_{2},\dots,c_{n_{\boldsymbol{U}}}]$ where only one of $c_{i},i=1\dots c_{n_{\boldsymbol{U}}}$ is $1$ and the rest are 00. Speech units are predicted at 50Hz.
@@ 4.1 Mesh-To-Speech
Regressing audio from 3D facial motion has, to our knowledge, not been done before. To narrow down the architecture search space, we take inspiration from a recent SVTS model, namely Choi et al. [choi2023lip2speechunit].
Architecture. The architecture consists of an input feature encoder, a conformer sequence encoder and two prediction heads, a speech unit classifier, and a mel spectrogram regressor. These two outputs contain enough information that a pretrained off-the-shelf vocoder [choi2023lip2speechunit] can turn them into output audio. Since the input is not a video, but a 3D animation, we replace the original lip video ResNet encoder that Choi et al. [choi2023lip2speechunit] use with an MLP that takes 3D lip vertex coordinates as the input. We keep the rest of the architecture the same. The mesh-to-speech architecture is depicted in Fig. 2 and can be written
@@
Supervision. We follow the same supervision scheme as Choi et al. [choi2023lip2speechunit]. The loss consists of two terms. The first term is an L1 loss between the input and output mel specrograms:
@@ 4.2 Speech-Driven Facial Animation Diffusion
Architecture. The core of THUNDER consists of a denoiser in the form of a transformer decoder [vaswani2017attention]. Similar to other contemporary 3D animation methods (e.g. Zhao et al. [zhao2024media2face] and Sun et al. [sun2024diffposetalk]), THUNDER processes a noisy sequence of expression parameters $\boldsymbol{\tilde{x}}_{(d)}^{1:T}$, where $d$ represents the diffusion timestep. These parameters serve as the query inputs to the transformer. The keys and values for the transformer are comprised of the input audio features $\mathbf{a}^{1:T}$ and a diffusion step embedding $\mathcal{T}(d)$, concatenated along the temporal dimension. The function $\mathcal{T}$ is defined as a sinusoidal timestep function similar to Ho et al. [ho2020denoising]. We do not incorporate additional animation-controlling conditions (such as style vectors [cudeiro2019capture, fan2022faceformer, xing2023codetalker], emotion vectors [danecek2023emote, Peng2023_EmoTalk] or CLIP embeddings [zhao2024media2face]) as this is not the goal of this work. However, the design can easily be augmented with additional input conditions. All inputs are projected into the transformer’s feature dimension, $f=128$, using dedicated learnable affine layers. Positional encoding within the transformer’s self-attention utilizes the ALiBi mechanism [press2021train]. The cross-attention layers use a diagonal binary mask to facilitate the attention mechanism, and no additional positional encoding. The model predicts the reconstructed expression parameters $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$. The transformation can be formally described by the following equation:
@@
[…]
@@
[…]
@@
$\mathcal{L}_{total}=w_{m2s}\mathcal{L}_{m2s}+\mathcal{L}_{rec}+\mathcal{L}_{vel},$ (8)
with $w_{mts}=1$. The model is trained using classifier-free guidance [ho2022cfg]. The audio condition is randomly dropped and replaced with a learnable vector with 20% probability.
Inference. At inference time, we initialize the first step by sampling the Gaussian noise: $\boldsymbol{\widehat{x}}_{(D)}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$. We then proceed to denoise for $D$ diffusion steps. In each step, we first employ the denoiser $\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\mathbf{a}^{1:T})$ to obtain $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$. Then, we compute $\boldsymbol{\tilde{x}}_{(d-1)}$ by adding the expected amount of noise back to obtain $\boldsymbol{\tilde{x}}_{(d-1)}$. Since THUNDER is trained with classifier-free guidance, it is possible to trade-off fidelity of the lip-sync for increased diversity by combining $s_{a}\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\mathbf{a}^{1:T})+(1-s_{a})\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\O)$ during the denoising process. However, we find our results are diverse enough without this and so we effectively set $s_{a}=1$.